LBC restart after summer shutdown:
UT 2011 SEP 08 & 09

Observer: D.Thompson, O.Kuhn, A.Rakich (Tucson), J.Hill (LBT)
Telescope Operator: G.Bechetti
Instrument Support: D.Officer
Telescope Support: J.Hill
SW support:

Summary

Reboot the CMU before observing starts
Does not seem to be necessary, top output looks OK with no excessive memory usage.

Check CMU system time vs our GPS:
[root@CMU root]# ntpdate -q 10.144.0.211
server 10.144.0.211, stratum 16, offset 82.083913, delay 0.02597
 8 Sep 02:04:18 ntpdate[13034]: no server suitable for synchronization found

Obtain biases and a dark:
Lights on in dome, so I only took 1 zero on each side. Look normal enough. Red bias on chip 2 is 481, on Blue it is 478.

Tried running LBC through the VPN from my office (DThompson). Initially the connection seemed fast, but when I tried displaying something on DS9 everything slowed waaay down to the point of unusability. As soon as the DS9 window was closed, other windows became responsive again. Moving to TO station for further work.

2:53 Talked to mtn. Andrew & JHill decided to start with bent Gregorian tests because we are out of thermal equil. Removing retroreflector and filling LBCs and PISCES.

10:52 Finally opening! Starting with DX RFBG.
11:12 Problems reading out acquire image from AzCam, John reports that it worked fine some hours ago. Switching to LUCI port.

11:24 Starlight on the ACQUIRE mode AGw frame! Badly aberrated. Adjusting M2 position

UT 2011 SEP 09

No data - clouds and occasional precipitation kept us closed all night.

-- DavidThompson - 08 Sep 2011
Topic revision: r3 - 09 Sep 2011, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback